How do you determine the units digit of 3 to the 99th power???
Answer by scott8148(6628)

3^1=3 ... 3^2=9 ... 3^3=27 ... 3^4=81 ... 3^5=243 (same units digit as 3^1)

the unit digits repeat in a pattern of four ... 99 divided by 4 is 24 with a remainder of 3

so the units digit for 3^99 is the third units digit in the pattern, which is 7